A grocery store needs to allocate overhead expenses of $309 across its three departments based on their sales. If refrigerated foods accounted for $748 in sales, non-refrigerated foods accounted for $907 in sales, and non-food products accounted for $612 in sales, what amount of overhead expenses should be allocated to non-food products?
Answer #1

Total overhead expenses = $309

sales ratio of 3 departments = 748 : 907 : 612

Total sales = 748+907+612= 2267.

Therefore Overhead expense allocated to non-food products = Total Overhead*(sales of Non-food products/Total sales)

=309 * (612/2267)

=$83.418